OpenStudy (anonymous):

What is the greatest common factor of 36x3 and 8x?

OpenStudy (anonymous):

The answer is 4x factors of 36x^3=2*2*3*3*x*x*x factors of 8x = 2*2*2*x common are GCF 2*2*x GCF = 4x

OpenStudy (anonymous):

thank u